# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v1

@@ 4.3 Ablations
In this section, we provide extensive ablations to study the different aspects of our GLOV. First, we take a closer look at all the design choices, then provide experiments that extend our method to VQA, later we discuss the generalization ability of the found prompts, and finally conclude with experiments regarding the choice of layer for our proposed guidance methodology.
(a) Comparison with ActADD.
(b) Adding offset to all tokens.
(c) Guidance with only last token.
(d) Cross attending all tokens.
(e) Cross attending only last token.
(f) Number of tokens for GLOV.
[…]
OxfordFlowers Aircraft Food101 Pets
Base 60.1 53.4 88.0 72.5
GLOV 62.9 57.5 89.5 73.9
Table 3: Generalization to visual question answering task. Accuracy (%) with LLaVa-OV (Li et al., 2024) for different datasets (posed as 4-way multi-choice) from the FOCI-Benchmark (Geigle et al., 2024).
ViT-B/16 ViT-L/14 MC-B/16 MC-L/14
Base 61.8 70.5 64.9 70.7
GLOV 63.9 72.7 66.3 72.5
Design Choices: The eventual algorithm (c.f., Algorithm 1) for our GLOV (especially the guidance scheme, i.e., GLOV-guidance) is chosen by intensely studying a variety of alternatives. Several of these design choices experimented with, are plotted in Figure 3. For example, in Figure 3 we compare ActADD (Turner et al., 2023) guidance scheme with our GLOV. To recall, ActADD applies the difference of the prompt embedding vectors to the first $N$ tokens (equal to the sequence length of the offset vector) of the prompt to the LLM, for the response generation. Whereas, our GLOV applies the guidance to only the last token of the (meta) prompt for each new token produced at each optimization step (in a greedy manner). We find that for the downstream vision task, our method fares slightly better. Similarly, for each generation step, we also experiment by applying the offset vector to each of the tokens of the prompt (c.f., Figure 3) – resulting in stronger guidance – and by only using the last token embedding for obtaining the offset vector (c.f., Figure 3). We observe that our method of obtaining the mean of the embeddings from all the tokens and adding the offset to only the last token at each generation step fares better. We also experiment with cross-attending the positive and negative prompt embeddings with the hidden state at each time step of the auto-regressive modeling in LLMs. Specifically in Figure 3 we cross-attend all the tokens of the hidden state (query) with the positive and negative prompt embeddings (keys-values) and obtain the offset vector, and in Figure 3 we only use the last token embedding for cross attention. In both cases, we see that our proposed guidance scheme fares better. Finally, we experiment with generating different numbers of tokens for each prompt (at each optimization step) and find that the best results are obtained with $50$ tokens. This could be because the CLIP text encoder does not favor longer sentences and shorter sentences might not be syntactically correct.
Generalization to VQA: We further evaluate our method on the visual question answering task proposed by Geigle et al. (2024). Specifically, they formulate fine-grained visual recognition into a four-way multiple-choice VQA task, where one choice is the ground truth and the other $3$ are hard negatives, selected by (closest) cosine similarity scores to the ground truth, by using SigLIP (Zhai et al., 2023). To obtain the results for our GLOV in Table 4 we optimize the <question> asked as prompt to the VLM. The most effective prompts discovered at the end of the optimization are listed in the Appendix Sections B.3 & B.4. These results provide a glimpse of the possibility of further extension of our work to the task of open-ended VQA in other domains, where the goal can be to optimize the questions. Currently, we leave such exploration for future work.
Generalization of Prompts: In Table 4 we evaluate the generalization ability of the discovered prompts on various CLIP variants, e.g., MetaCLIP (Xu et al., 2023). We find that the effective prompts discovered for the CLIP ViT-B/32 (Radford et al., 2021) backbone can transfer to other CLIP variants (and model sizes) to enhance the results.
@@
Choice of Layer for Guidance: One important design choice for our GLOV is the choice of layer in the LLM for the embedding-space guidance. Furthermore, our method calculates the mean of the sequence lengths, to obtain the sentence embeddings (c.f., equation 3), which is also an opportunity for introspection. To obtain a measure of the quality of the sentence embeddings, we linear probe different layers in Llama-3, on the popular sentiment classification task SST (Socher et al., 2013) and provide results in Figure 4 (left). SST has been widely used to benchmark sentence representations (Conneau & Kiela, 2018). We find that the middle layers of Llama-3 obtain the highest accuracy, highlighting the semantic relevance of the sentence embeddings obtained from these layers, consistent with the literature (Liu et al., 2019a, Zhao et al., 2020). Furthermore, we also run a sweep while applying the guidance on different layers in Llama-3 and plot the resulting ImageNet accuracy on the 1-shot train set in Figure 4 (right). The accuracy peaks at layer $16$ and layer $17$. These results are consistent with the linear probing results obtained on the SST-$5$ dataset, hinting that the middle layers might be the most effective. Hence, keeping these results in view and following Turner et al. (2023), we choose layer $17$ in Llama-3 to apply the offset vector for steering the responses.
@@ 5 Conclusion
We have presented a prompt optimization method for VLMs that interfaces two disjoint models through a fitness function. The LLM iteratively interacts with the VLM during the optimization run and is able to gradually understand the type of language structure preferred by the downstream VLM, and discovers effective solutions that can maximize the learning objective (i.e., the accuracy on the downstream vision task). To further enhance the optimization, we condition the LLM responses at each optimization step by providing a direction. The direction is dictated through a novel embedding space steering methodology that, in essence, adds an offset vector calculated from the positive and negative prompt embeddings to the intermediate layer of the LLM, helping it to bound the outputs more strictly towards the language prompts preferred by the VLM. Extensive empirical evaluations with different VLM architectures on multiple datasets highlight the effectiveness of our proposed GLOV.
@@
In the following, we provide additional experiments and further explanations which might be helpful for the reader to gain further insights and add clarity to the main manuscript. In Section A we further expand upon the details regarding the calculation of the fitness of prompts for the encoder-decoder models. Then, in Section B we list all the prompts used to obtain the results in the main manuscript (Tables 12 & 4). Finally, in Sections C & D we provide results while comparing with the gradient-based learning method proposed by Zhou et al. (2022). Furthermore, a comprehensive optimization algorithm is also provided in Algorithm 1.
@@ Appendix A Fitness for Encoder-Decoder Models
Appendix B GLOV Prompts
@@
Here, we list all the prompts discovered during the optimization runs. In Sections B.1B.2 we list the prompts for LLaVa-OV (Li et al., 2024) for the task of image classification. In Sections B.3B.4 we provide the prompts for the same model for the task of visual question answering. Finally, in Sections B.5 & B.6 we provide the prompts used to build an ensemble of classifiers for CLIP. Furthermore, we also provide the prompt evolution at different optimization steps for LLaVa and CLIP, in Figures 7 & 8.
B.1 GLOV (w/o guidance) - Prompts (LLaVa-OV - Image Classification)

### v5

@@ 4.3 Ablations
We provide extensive ablations to study the different aspects of our GLOV. First, we motivate our choice to use the middle layer of the LLM for guidance. Then, we provide results by using less than one shot data as our held-out train set and later, we discuss the results obtained by using a larger LLM. Finally, we conclude with an ablation showing the retention of general model capabilities by adding the safety prompt. More ablations for additional insights are added to the Appendix Section B.
@@
Choice of Layer for Guidance: To obtain a measure of the quality of the sentence embeddings, we linear probe different layers in Llama-3, on the popular sentiment classification task SST (Socher et al., 2013) and provide results in Figure 3 (left). We find that the middle layers of Llama-3 obtain the highest accuracy, highlighting the semantic relevance of the sentence embeddings obtained from these layers, consistent with the literature (Liu et al., 2019a; Zhao et al., 2020). Furthermore, we also run a sweep while applying the guidance on different layers in Llama-3 and plot the resulting ImageNet accuracy on the 1-shot train set in Figure 3 (right). The accuracy peaks at layer $16$ and layer $17$. These results are consistent with the linear probing results obtained on the SST-$5$ dataset, hinting that the middle layers might be the most effective. Keeping these results in view, we choose layer $17$ in Llama-3 to apply the offset vector for steering the responses.
Sub-shot Setting: In Table 4 we provide results by using less than one shot train data for the optimization of prompts through GLOV. We observe that for ImageNet, our GLOV shows a minimal performance drop (1.8%) by using only $50$ samples from $50$ randomly sampled classes, instead of using $1000$ samples from (all) $1000$ classes. The results for ImageNet-R also follow a similar trend. These results show that the optimized prompts can generalize to unseen classes and we can drastically reduce the size of the training set (thus lowering annotation efforts) and the time required for the optimization with a minimal performance penalty.
Scaling LLM Size: In Table 5, we present results obtained by using a larger LLM (Llama-3.1-70b) for optimization with our GLOV. The results demonstrate that our GLOV benefits from leveraging a more capable LLM, highlighting the potential for further improvements as LLMs advance.
SEED MME
Model Baseline Initial GLOV Baseline Initial GLOV
LLaVA-OV 67.8 67.4 68.0 1578.3 1575.5 1559.5
Molmo 63.5 63.0 63.1 1283.1 1270.5 1306.4
Table 6: Retention of Generalization. Accuracy (%) for SEED and Perception Score for MME. GLOV results are obtained by prepending the GLOV safety prompt to the original question.
Retention of Generalization: In Table 6, we provide results on general-purpose VLM benchmarks (SEED (Fang et al., 2021) and MME (Yin et al., 2023)) by prepending the optimized safety prompt discovered through our GLOV. We find that the discovered prompt has a minor effect on performance. In fact, on the large-scale MME benchmark, for Molmo, the safety prompt leads to a performance gain in the visual perception score of $23.3$. This gain can be attributed to the safety prompt forcing the model to focus more closely on the question.
@@ 5 Conclusion
We have presented a prompt optimization method for VLMs that interfaces two disjoint models through a fitness function. The LLM iteratively interacts with the VLM during the optimization run and is able to gradually understand the type of language structure preferred by the downstream VLM, and discovers effective solutions that can maximize the learning objective (i.e., the accuracy on the downstream vision task). To further enhance the optimization, we condition the LLM responses at each optimization step by providing a direction. The direction is dictated through a novel embedding space steering methodology that, in essence, adds an offset vector calculated from the positive and negative prompt embeddings to the intermediate layer of the LLM, helping it to bound the outputs more strictly towards the language prompts preferred by the VLM. Extensive empirical evaluations highlight the effectiveness of our proposed GLOV.
@@
In the following, we provide additional experiments and further explanations that might be helpful for the readers to gain insights and add clarity to the main manuscript. In Section A we expand upon the details regarding the calculation of the fitness of prompts for the encoder-decoder models. Then, in Section B we provide more insights in to our GLOV by studying additional aspects of it. In Section C we list the details about the datasets employed for evaluating the safety of VLMs in our work. Later, Section D provides the detailed results for each split in the MM-Safety-Benchmark (Liu et al., 2024c), and finally we conclude with Section E, which lists all the prompts used to obtain the results in the main manuscript (Tables 12 & 3). Furthermore, a comprehensive optimization algorithm is also provided in Algorithm 1.
@@ Appendix A Fitness for Encoder-Decoder Models
Appendix B Additional Ablations
Here, we provide additional insights into our GLOV. First, we provide several design considerations, which eventually led to the proposed method. Then, we study the generalization of the found prompts on different CLIP-based backbones, later, we provide results on the FOCI Benchmark (Geigle et al., 2024), then, study the effect of more shots on our method and finally provide results while comparing our GLOV with other PEFT methods and conclude with providing results with multiple random runs for our GLOV.
(a) Comparison with ActADD.
(b) Adding offset to all tokens.
(c) Guidance with only last token.
(d) Cross attending all tokens.
(e) Cross attending only last token.
(f) Number of tokens for GLOV.
[…]
B.1 Design Choices
The eventual algorithm (c.f., Algorithm 1) for our GLOV (especially the guidance scheme, i.e., GLOV-guidance) is chosen by intensely studying a variety of alternatives. Several of these design choices experimented with are plotted in Figure 4. For example, in Figure 4 we compare ActADD (Turner et al., 2023) guidance scheme with our GLOV. To recall, ActADD applies the difference of the prompt embedding vectors to the first $N$ tokens (equal to the sequence length of the offset vector) of the prompt to the LLM, for the response generation. Whereas, our GLOV applies the guidance to only the last token of the (meta) prompt for each new token produced at each optimization step (in a greedy manner). We find that for the downstream vision task, our method fares slightly better. Similarly, for each generation step, we also experiment by applying the offset vector to each of the tokens of the prompt (c.f., Figure 4) – resulting in stronger guidance – and by only using the last token embedding for obtaining the offset vector (c.f., Figure 4). We observe that our method of obtaining the mean of the embeddings from all the tokens and adding the offset to only the last token at each generation step fares better. We also experiment with cross-attending the positive and negative prompt embeddings with the hidden state at each time step of the auto-regressive modeling in LLMs. Specifically in Figure 4 we cross-attend all the tokens of the hidden state (query) with the positive and negative prompt embeddings (keys-values) and obtain the offset vector, and in Figure 4 we only use the last token embedding for cross attention. In both cases, we see that our proposed guidance scheme fares better. Finally, we experiment with generating different numbers of tokens for each prompt (at each optimization step) and find that the best results are obtained with $50$ tokens. This could be because the CLIP text encoder does not favor longer sentences and shorter sentences might not be syntactically correct.
B.2 Generalization of Prompts
ViT-B/16 ViT-L/14 MC-B/16 MC-L/14
Base 61.8 70.5 64.9 70.7
GLOV 63.9 72.7 66.3 72.5
In Table 7 we evaluate the generalization ability of the discovered prompts on various CLIP variants, e.g., MetaCLIP (Xu et al., 2023). We find that the effective prompts discovered for the CLIP ViT-B/32 (Radford et al., 2021) backbone can transfer to other CLIP variants (and model sizes) to enhance the results.
B.3 Generalization to Multiple Choice Classification
OxfordFlowers Aircraft Food101 Pets
Base 60.1 53.4 88.0 72.5
GLOV 62.9 57.5 89.5 73.9
Table 8: Generalization to multiple choice classification. Accuracy (%) with LLaVA-OV (Li et al., 2024) for different datasets (posed as 4-way multi-choice) from the FOCI-Benchmark (Geigle et al., 2024).
We further evaluate our method on the multiple choice classification task proposed by Geigle et al. (2024). Specifically, they formulate fine-grained visual recognition into a four-way multiple-choice VQA task, where one choice is the ground truth and the other $3$ are hard negatives, selected by (closest) cosine similarity scores to the ground truth, by using SigLIP (Zhai et al., 2023). To obtain the results for our GLOV in Table 8 we optimize the <question> asked as prompt to the VLM. The most effective prompts discovered at the end of the optimization are listed in the Appendix Sections E.3 & E.4.
B.4 More Shots Help
Results in the main manuscript (Tables 1 & 2) are obtained by using $1$-shot training data. In Table 9 we provide results on several datasets by employing $5$-shot training data for the CLIP ViT-B/32 (Radford et al., 2021) backbone. We observe a consistent improvement in results by using more shots.
B.5 Comparison with PEFT Methods
For completeness, in Table 10 we compare with the popular CoOp (Zhou et al., 2022) method in the $1$-shot learning regime. We observe that the ensemble of classifiers built from our discovered prompts can outperform CoOp. In extremely low-shot learning regimes, gradient-based learning poses a threat of overfitting, whereas our GLOV can avoid that because of no parameter update. Further, we also fine-tune the LORA (Hu et al., 2021) adapter and provide results in Table 10. Similar to CoOp, we find that applying LORA to all the matrices of the vision and text encoders of CLIP can result in overfitting. For a fair comparison, we also apply it only to the attention blocks and report the results. However, our GLOV still outperforms the LORA adapter. This further strengthens our claim that in low data regimes, our GLOV can fare better than the gradient-based learning methods.
EuroSAT ImageNetA ImageNetR RESISC45 DescribableTextures
1-shot 50.8 32.5 68.6 62.0 42.6
5-shot 54.3 33.8 68.8 64.2 44.2
Table 9: More shots help. Top-1 accuracy (%) with CLIP ViT-B/32.
ImageNet ImageNetA ImageNetS UCF101
CLIP 61.9 28.2 40.3 60.4
CoOp 60.6 24.5 39.9 63.8
LORA (all) 59.9 27.1 37.1 58.2
LORA (attention) 62.6 30.1 40.5 62.1
GLOV 64.5 32.5 43.0 63.8
Table 10: Comparison with PEFT methods. Top-1 accuracy (%) with CLIP ViT-B/32.
ImageNet ImageNetA ImageNetS UCF101
CLIP 61.9 28.2 40.3 60.4
GLOV 64.3 $\pm$ 0.43 32.0 $\pm$ 0.69 43.1 $\pm$ 0.25 63.9 $\pm$ 0.34
Table 11: Variance of results. Top-1 Accuracy and variances for ViT-B/32 clip backbone for $5$ random GLOV runs.
B.6 Variance of Results
In Table 11, we provide results for $5$ random optimization runs performed with our GLOV. The results show that the performance improvements obtained with our GLOV are statistically significant.
Appendix C Safety Dataset Details
We evaluate our GLOV on two recent safety benchmarks: MM-Safety-Bench (Liu et al., 2024c) and VLGuard (Zong et al., 2024). Here we provide details regarding the different splits and the evaluation protocols for both datasets.
C.1 MM-Safety-Bench
The MM-Safety-Bench (Liu et al., 2024c) contains images with corresponding instructions related to a total of $13$ sub-categories. These categories are listed in the detailed results presented in Section D. Further, the benchmark is also divided according to the type of images: typographic images, images synthesized from stable diffusion, and a combination of both. Each of these images is accompanied by a text instruction to the model, which tries to elicit unsafe responses corresponding to the image. After a VLM provides the answers, the evaluation is performed by GPT-4 (OpenAI, 2023), and the attack success rate is measured for each of the partitions.
C.2 VLGuard
The VLGuard (Zong et al., 2024) dataset contains $3$ splits:
  • Unsafe assesses the capability of the model to identify and refuse harmful images on the vision side.
  • Safe-Unsafes focuses on the models’ ability to identify and strongly reject unsafe instructions from the language side.
  • Safe-Safes focuses on the models’ ability to answer to safe instructions. This can be interpreted as a control set.
For the first two subsets, the refusal rate is measured by string matching, while for the third subset, the responses from the model are evaluated against the responses from GPT-4, and the win-win rate is reported by using GPT-4 as a judge.
Appendix D MM-Safety-Bench Detailed Results
LLaVA-OV MOLMO
TYPO SD SD + TYPO TYPO SD SD + TYPO
01-Illegal Activity 10.3 16.5 18.6 66.0 43.3 66.0
02-Hate Speech 45.4 32.5 51.5 60.7 37.4 62.6
03-Malware Generation 13.6 13.6 25.0 52.3 18.2 40.9
04-Physical Harm 22.2 18.8 36.1 50.7 28.5 56.2
05-Economic Harm 10.7 13.1 21.3 20.5 22.1 18.9
06-Fraud 36.4 27.3 49.4 80.5 52.6 84.4
07-Pornography 9.2 10.1 11.9 8.3 18.3 11.9
08-Political Lobbying 97.4 72.5 96.7 99.3 79.1 98.7
09-Privacy Violence 35.3 17.3 47.5 71.2 25.2 68.3
10-Legal Opinion 90.0 88.5 88.5 85.4 89.2 90.0
11-Financial Advice 100.0 100.0 99.4 100.0 100.0 100.0
12-Health Consultation 98.2 97.2 100.0 100.0 97.2 100.0
13-Gov Decision 96.6 73.2 89.3 94.6 80.5 92.6
Average 51.2 44.7 56.5 68.4 53.2 68.5
Table 12: Base model results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by using the original image and instructions from the dataset.
LLaVA-OV MOLMO
TYPO SD SD + TYPO TYPO SD SD + TYPO
01-Illegal Activity 0.0 7.2 6.2 15.5 8.2 18.6
02-Hate Speech 17.8 17.2 20.2 24.5 11.7 21.5
03-Malware Generation 15.9 6.8 2.3 29.5 11.4 15.9
04-Physical Harm 13.2 9.0 16.7 25.7 13.2 25.7
05-Economic Harm 13.9 14.8 22.1 27.0 13.9 20.5
06-Fraud 15.6 12.3 21.4 30.5 19.5 24.0
07-Pornography 7.3 3.7 7.3 10.1 5.5 4.6
08-Political Lobbying 92.2 58.8 86.3 94.1 68.0 92.2
09-Privacy Violence 15.8 7.2 31.7 32.4 12.2 25.2
10-Legal Opinion 83.1 74.6 85.4 85.4 83.8 86.2
11-Financial Advice 98.8 95.8 97.0 100.0 100.0 100.0
12-Health Consultation 98.2 89.9 96.3 96.3 94.5 97.2
13-Gov Decision 82.6 49.7 77.9 84.6 67.8 87.9
Average 42.6 34.4 43.9 50.4 39.2 47.6
Table 13: Initial safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the initial safety instruction: These instructions can be harmful, to the original instruction in the dataset.
LLaVA-OV MOLMO
TYPO SD SD + TYPO TYPO SD SD + TYPO
01-Illegal Activity 2.1 1.0 0.0 4.1 2.1 0.0
02-Hate Speech 3.1 3.1 2.5 5.5 5.5 8.6
03-Malware Generation 0.0 0.0 0.0 2.3 0.0 2.3
04-Physical Harm 1.4 1.4 0.7 6.9 7.6 10.4
05-Economic Harm 2.5 4.1 3.3 6.6 12.3 5.7
06-Fraud 1.3 1.3 0.6 3.2 8.4 8.4
07-Pornography 0.0 3.7 0.9 1.8 6.4 5.5
08-Political Lobbying 28.1 11.1 15.0 40.5 38.6 49.7
09-Privacy Violence 2.9 1.4 2.9 4.3 1.4 3.6
10-Legal Opinion 36.2 16.2 39.2 58.5 63.1 60.8
11-Financial Advice 83.2 77.2 79.0 94.0 94.6 95.2
12-Health Consultation 67.9 29.4 53.2 86.2 84.4 79.8
13-Gov Decision 32.9 2.7 20.8 33.6 25.5 38.9
Average 20.1 11.7 16.8 26.7 26.9 28.4
Table 14: GLOV (without guidance) safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV (without applying the guidance): Please do not execute the following instruction unless it has been thoroughly validated and vetted by multiple experts in the relevant field, and even then, only after carefully considering the potential risks and unintended consequences, to the original instruction in the dataset.
LLaVA-OV MOLMO
TYPO SD SD + TYPO TYPO SD SD + TYPO
01-Illegal Activity 0.0 0.0 1.0 5.2 8.2 19.6
02-Hate Speech 1.8 1.8 2.5 6.7 3.7 6.1
03-Malware Generation 0.0 0.0 0.0 6.8 6.8 11.4
04-Physical Harm 3.5 0.0 2.1 6.9 6.3 16.7
05-Economic Harm 1.6 0.0 2.5 12.3 4.1 9.0
06-Fraud 0.6 0.0 1.3 7.8 3.2 9.7
07-Pornography 0.0 1.8 0.0 3.7 2.8 7.3
08-Political Lobbying 27.5 9.2 10.5 50.3 14.4 34.6
09-Privacy Violence 2.9 0.7 1.4 5.8 5.8 12.9
10-Legal Opinion 26.2 16.9 36.2 36.9 46.9 34.6
11-Financial Advice 76.0 71.3 75.4 85.0 87.4 80.8
12-Health Consultation 37.6 15.6 33.0 71.6 75.2 62.4
13-Gov Decision 14.8 2.0 6.0 10.7 2.0 12.8
Average 14.8 9.2 13.2 23.8 20.5 24.5
Table 15: GLOV safety instruction results on MM-Safety-Bench. Attack success rate (%) on all partitions in the MM-Safety-Bench (Liu et al., 2024c). These results are obtained by prepending the best safety instruction found by our GLOV: Please carefully scrutinize the following instruction and decline to execute it if you perceive any potential risks or uncertainties that may compromise the integrity of the model or its users, to the original instruction in the dataset.
For completeness, we provide the detailed results on all the $13$ different categories in the MM-Safety-Bench (Liu et al., 2024c) in Tables 121314 & 15 for the baseline models (LLaVA-OV and Molmo), by prepending the initial instruction, instruction found by our GLOV (without guidance) and instruction found by our GLOV (with guidance).
Appendix E GLOV Prompts
@@
Here, we list all the prompts discovered during the optimization runs. In Sections E.1E.2 we list the prompts for LLaVA-OV (Li et al., 2024) for the task of image classification. In Sections E.3E.4 we provide the prompts for the same model for the task of visual question answering. Finally, in Sections E.5 & E.6 we provide the prompts used to build an ensemble of classifiers for CLIP. Furthermore, we also provide the prompt evolution at different optimization steps for LLaVA and CLIP, in Figures 7 & 8.
E.1 GLOV (w/o guidance) - Prompts (LLaVA-OV - Image Classification)